# Graph-level Representation Learning with Joint-Embedding Predictive Architectures

### v1

@@ Abstract
Joint-Embedding Predictive Architectures (JEPAs) have recently emerged as a novel and powerful technique for self-supervised representation learning. They aim to learn an energy-based model by predicting the latent representation of a target signal $y$ from a context signal $x$. JEPAs bypass the need for data augmentation and negative samples, which are typically required by contrastive learning, while avoiding the overfitting issues associated with generative-based pretraining. In this paper, we show that graph-level representations can be effectively modeled using this paradigm and propose Graph-JEPA, the first JEPA for the graph domain. In particular, we employ masked modeling to learn embeddings for different subgraphs of the input graph. To endow the representations with the implicit hierarchy that is often present in graph-level concepts, we devise an alternative training objective that consists of predicting the coordinates of the encoded subgraphs on the unit hyperbola in the 2D plane. Extensive validation shows that Graph-JEPA can learn representations that are expressive and competitive in both graph classification and regression problems.
@@ 1 Introduction
Graph data is ubiquitous in the real world due to its ability to universally abstract various concepts and entities (Ma & Tang, 2021; Veličković, 2023). To deal with this peculiar data structure, Graph Neural Networks (GNNs) (Scarselli et al., 2008; Kipf & Welling, 2016a; Gilmer et al., 2017; Veličković et al., 2017) have established themselves as a staple solution. Nevertheless, most applications of GNNs usually rely on ground-truth labels for training. The growing amount of graph data in fields such as bioinformatics, chemoinformatics, and social networks has made manual labeling laborious, sparking significant interest in unsupervised graph representation learning.
@@
A particularly emergent area in this line of research is self-supervised learning (SSL). In SSL, alternative forms of supervision are created stemming from the input signal. This process is then typically followed by invariance-based or generative-based pretraining (Liu et al., 2023; Assran et al., 2023). Invariance-based approaches optimize the model to produce comparable embeddings for different views of the input signal. A common term associated with this procedure is contrastive learning (Tian et al., 2020). Typically, these alternative views are created by a data augmentation procedure. The views are then passed through their respective encoder networks (which may share weights), as shown in Figure 1a. Finally, an energy function, usually framed as a distance, acts on the latent embeddings. In the graph domain, several works have applied contrastive learning by designing graph-specific augmentations (You et al., 2020), using multi-view learning (Hassani & Khasahmadi, 2020) and even adversarial learning (Suresh et al., 2021). This pretraining strategy is effective but comes with several drawbacks i.e., the necessity to augment the data and process negative samples. In order to learn embeddings that are useful for downstream tasks, the augmentations must also be non-trivial.
@@
Generative methods on the other hand typically remove or corrupt portions of the input and predict them using an autoencoding procedure (Vincent et al., 2010; He et al., 2022), or rely on autoregressive modeling (Brown et al., 2020; Hu et al., 2020). Figure 1b depicts the typical instantiation of these methods: The input signal $x$ is fed into an encoder network that constructs the latent representation and a decoder generates $\hat{y}$, the data corresponding to the target signal $y$. The energy function is then applied in data space, often as a reconstruction term (Bengio et al., 2013). Generative models generally display strong overfitting tendencies since they have to estimate the data distribution (implicitly or explicitly), so the latent representations must be directly descriptive of the whole data space. This can be very problematic for graphs given that they live in a non-Euclidean and inhomogenous data space. Nevertheless, masked autoencoding has recently shown promising results in the graph domain with appropriately designed models (Hou et al., 2022; Tan et al., 2023).
@@
Figure 1: Illustration of the three main SSL approaches: (a) Joint-Embedding Architectures learn to create similar embeddings for inputs x and y that are compatible with each other and dissimilar embeddings for inputs that are not compatible. This compatibility is implemented in practice by creating different views of the input data. (b) Generative Architectures reconstruct a signal $y$ from a compatible signal $x$ by conditioning the decoder network on additional (potentially latent) variables $z$. (c) Joint-Embedding Predictive Architectures act as a bridge: They utilize a predictor network that processes the context $x$ and is conditioned on additional (potentially latent) factors, to predict the embedding of the target $y$ in latent space.
@@
Inspired by the innovative Joint-Embedding Predictive Architecture (JEPA) (LeCun, 2022; Assran et al., 2023), we propose Graph-JEPA to learn graph-level representations by bridging contrastive and generative models. As illustrated in Figure 1c, a JEPA has two encoder networks that receive the input signals and produce the corresponding representations. Notably, the two encoders can be different models and don’t need to share weights. A predictor module outputs a prediction of the latent representation of one signal based on the other, possibly conditioned on another variable. Graph-JEPA does not require any negative samples or data augmentation and by operating in the latent space it avoids the pitfalls associated with learning high-level details needed to fit the data distribution. However, the graph domain presents us with additional challenges, namely: context and target extraction; designing a latent prediction task that is optimal for graph-level concepts; and learning expressive representations. In response to these questions, we design Graph-JEPA with a specific masked modeling objective. The input graph is first divided into several subgraphs, and then the latent representation of randomly chosen target subgraphs is predicted given a context subgraph. The subgraph representations are consequently pooled to create a graph-level representation.
@@
The nature of graph-level concepts is often assumed to be hierarchical (Ying et al., 2018), thus we conjecture that the typical latent reconstruction objective used in current JEPA formulations is not enough to provide optimal downstream performance. To this end, we design a prediction objective that starts by expressing the target subgraph encoding as a high-dimensional description of the hyperbolic angle. The predictor module is then tasked with predicting the location of the target in the 2D unit hyperbola. This prediction is compared with the target coordinates, obtained by using the aforementioned hyperbolic angle. Graph-JEPA outperforms popular contrastive and generative graph-level SSL methods on different datasets, while maintaining efficiency and ease of training. Notably, we observe from our experiments that Graph-JEPA can run up to 1.45x faster than Graph-MAE (Hou et al., 2022) and 8x faster than MVGRL (Hassani & Khasahmadi, 2020). Finally, we empirically demonstrate Graph-JEPA’s ability to learn highly expressive graph representations, showing it almost perfectly distinguishes pairs of non-isomorphic graphs that the 1-WL test cannot differentiate.
@@ 2.1 Self-Supervised Graph Representation Learning
Graph Neural Networks (GNNs) (Scarselli et al., 2008; Kipf & Welling, 2016a; Veličković et al., 2017; Hamilton et al., 2017; Xu et al., 2018; Wu et al., 2019) are now established solutions to different graph machine learning problems such as node classification, link prediction, and graph classification. Nevertheless, the cost of labeling graph data is quite high given the immense variability of graph types and the information they can represent. To alleviate this problem, SSL on graphs has become a research frontier. SSL methods on graphs can be divided into two major groups (Liu et al., 2023):
@@
Contrastive Methods. Contrastive learning methods usually consist of minimizing an energy function (Hinton, 2002; Gutmann & Hyvärinen, 2010) between different views of the same data. InfoGraph (Sun et al., 2019) maximizes the mutual information between the graph-level representation and the representations of substructures at different scales. GraphCL (You et al., 2020) works similarly to distance-based contrastive methods in the imaging domain. The authors first propose four types of graph augmentations and then perform contrastive learning based on them. The work of Hassani & Khasahmadi (2020) goes one step further by contrasting structural views of graphs. They also show that a large number of views or multiscale training does not seem to be beneficial, contrary to the image domain. Another popular research direction for contrastive methods is learning graph augmentations (Suresh et al., 2021) and AutoSSL (Jin et al., 2021), where the goal is to learn how to automatically leverage multiple pretext tasks effectively. Contrastive learning methods typically require a lot of memory due to data augmentation and negative samples. Graph-JEPA is much more efficient than typical formulations of these architectures given that it does not require any augmentations or negative samples. Another major difference is that the prediction in latent space in JEPAs is done through a separate predictor network, rather than using the common Siamese structure (Bromley et al., 1993)(Figure 1a vs c).
@@
Generative Methods. The goal of generative models is to recover the data distribution, an objective that is typically implemented through a reconstruction process. In the graph domain, most generative architectures that are also used for SSL are extensions of Auto-Encoder (AE) (Hinton & Zemel, 1993) and Variational Auto-Encoder (VAE) (Kingma & Welling, 2013) architectures. These models learn an embedding from the input data and then use a reconstruction objective with (optional) regularization in order to maximize the data evidence. Kipf & Welling (2016b) extended the framework of AEs and VAEs to graphs by using a GNN as an encoder and the reconstruction of the adjacency matrix as a training objective. However, the results on node and graph classification benchmarks with these embeddings are often unsatisfactory compared with contrastive learning methods, a tendency also observed in other domains (Liu et al., 2023). A recent and promising direction is masked autoencoding (MAE) (He et al., 2022), which has proved to be a very successful framework for the image and text domains. GraphMAE (Hou et al., 2022) is an instantiation of MAEs in the graph domain, where the node attributes are perturbed and then reconstructed, providing a paradigm shift from the structure learning objective of GAEs. S2GAE (Tan et al., 2023) is one of the latest GAEs, which focuses on reconstructing the topological structure but adds several auxiliary objectives and additional designs. Our architecture differs from generative models in that it learns to predict directly in the latent space, thereby bypassing the necessity of remembering and overfitting high-level details that help maximize the data evidence (Figure 1b vs c).
@@ 2.2 Joint-Embedding Predictive Architectures
Joint-Embedding Predictive Architectures (LeCun, 2022) are a recently proposed design for SSL. The idea is similar to both generative and contrastive approaches, yet JEPAs are non-generative since they cannot directly predict $y$ from $x$, as shown in Figure 1c. The energy of a JEPA is given by the prediction error in the embedding space, not the input space. These models can therefore be understood as a way to capture abstract dependencies between $x$ and $y$, potentially given another latent variable $z$. It is worth noting that the different models comprising the architecture may differ in terms of structure and parameters. An in-depth explanation of Joint-Embedding Predictive Architectures and their connections to human representation learning is provided by LeCun (2022). The most well-known works to use particular instantiations of JEPAs before the term was officially coined are BYOL (Grill et al., 2020) and SimSiam (Chen & He, 2021). There have recently been two works that implement JEPAs in practice, both in the vision domain: I-JEPA (Assran et al., 2023) and MC-JEPA (Bardes et al., 2023). I-JEPA provides an instantiation in the context of images, where the goal is to predict the latent embeddings of different image patches given a context patch. MC-JEPA instantiates the architecture in the context of learning optical flow and content features within a shared encoder. In this work, we propose the first JEPA for the graph domain and use it to learn graph-level representations in a self-supervised manner.
@@ 3 Method
Preliminary. Let a graph $G$ be defined as $G=(V,E)$ where $V=\{v_{1}\ldots v_{N}\}$ is the set of nodes, with a cardinality $|V|=N$, and $E=\{e_{1}\ldots e_{M}\}$ is the set of edges, with a cardinality $|E|=M$. We consider symmetric, unweighted graphs, although our method can be generalized to directed or weighted graphs. In this setting, $G$ can be represented by an adjacency matrix $A\in\{0,1\}^{N\times N}$, with $A_{ij}=1$ if nodes $v_{i}$ and $v_{j}$ are connected and $A_{ij}=0$ otherwise.
A general overview. Figure 2 provides an overview of the proposed architecture. The high-level idea of Graph-JEPA is to first divide the input graph into subgraphs (patches) (He et al., 2023) and then predict the representation of a randomly chosen target subgraph from a context subgraph (Assran et al., 2023). We would like to stress that this masked modeling objective is realized in latent space, without the need for augmentations or negative samples. The subgraph representations are then pooled to create a vectorized graph-level representation. Therefore Graph-JEPA can be described through a sequence of operations, namely: 1. Spatial Partitioning; 2. Subgraph Embedding; 3. Context and Target Encoding; 4. Latent Target Prediction.
Figure 2: An overview of Graph-JEPA. We first extract non-overlapping subgraphs (patches) (a.), perform a 1-hop neighborhood expansion (b.), and encode the subgraphs with a GNN (c.). After the subgraph encoding, one is randomly picked as the context and $m$ others as the targets (d.) and they are fed into their respective encoders (e.). The embeddings generated from the target encoder are used to produce the target subgraphs’ coordinates $\psi_{y}$. Finally, the predictor network is tasked with directly predicting the coordinates $\hat{\psi}_{y}$ for each target subgraph based on the context embedding and the positional embedding of each target subgraph (f.). A regression loss acts as the energy function $D$ between the predicted and target coordinates.

### v2

@@ Abstract
Joint-Embedding Predictive Architectures (JEPAs) have recently emerged as a novel and powerful technique for self-supervised representation learning. They aim to learn an energy-based model by predicting the latent representation of a target signal $y$ from the latent representation of a context signal $x$. JEPAs bypass the need for negative and positive samples, traditionally required by contrastive learning while avoiding the overfitting issues associated with generative pretraining. In this paper, we show that graph-level representations can be effectively modeled using this paradigm by proposing a Graph Joint-Embedding Predictive Architecture (Graph-JEPA). In particular, we employ masked modeling and focus on predicting the latent representations of masked subgraphs starting from the latent representation of a context subgraph. To endow the representations with the implicit hierarchy that is often present in graph-level concepts, we devise an alternative prediction objective that consists of predicting the coordinates of the encoded subgraphs on the unit hyperbola in the 2D plane. Through multiple experimental evaluations, we show that Graph-JEPA can learn highly semantic and expressive representations, as shown by the downstream performance in graph classification, regression, and distinguishing non-isomorphic graphs. The code will be made available upon acceptance.
@@ 1 Introduction
Graph data is ubiquitous in the real world due to its ability to universally abstract various concepts and problems (Ma & Tang, 2021; Veličković, 2023). To deal with this widespread data structure, Graph Neural Networks (GNNs) (Scarselli et al., 2008; Kipf & Welling, 2016a; Gilmer et al., 2017; Veličković et al., 2017) have established themselves as a staple solution. Nevertheless, most applications of GNNs usually rely on ground-truth labels for training. The growing amount of graph data in fields such as bioinformatics, chemoinformatics, and social networks makes manual labeling laborious, sparking significant interest in unsupervised graph representation learning.
@@
A particularly emergent area in this line of research is self-supervised learning (SSL). In SSL, alternative forms of supervision are created stemming from the input signal. This process is then typically followed by invariance-based or generative-based pretraining (Liu et al., 2023; Assran et al., 2023). Invariance-based approaches optimize the model to produce comparable embeddings for different views of the input signal. A common paradigm associated with this procedure is contrastive learning (Tian et al., 2020). Typically, these alternative views are created by a data augmentation procedure. The views are then passed through their respective encoder networks (which may share weights), as shown in Fig. 1a. Finally, an energy function, usually framed as a distance, acts on the latent embeddings. In the graph domain, several works have applied contrastive learning by designing graph-specific augmentations (You et al., 2020), using multi-view learning (Hassani & Khasahmadi, 2020) and even adversarial learning (Suresh et al., 2021). Invariance-based pretraining is effective but comes with several drawbacks i.e., the necessity to augment the data and process negative samples, which limits computational efficiency. In order to learn semantic embeddings that are useful for downstream tasks, the augmentations must also be non-trivial.
@@
Generative-based pretraining methods, on the other hand, typically remove or corrupt portions of the input and predict them using an autoencoding procedure (Vincent et al., 2010; He et al., 2022) or rely on autoregressive modeling (Brown et al., 2020; Hu et al., 2020). Fig. 1b depicts the typical instantiation of these methods: The input signal $x$ is fed into an encoder network that constructs the latent representation, and from it a decoder generates $\hat{y}$, the data corresponding to the target signal $y$. The energy function is then applied in data space, often through a reconstruction error. Generative models generally display strong overfitting tendencies van den Burg & Williams (2021) and can be non-trivial to train due to issues such as mode collapse Adiga et al. (2018). Moreover, the features they learn are not always useful for downstream tasks Meng et al. (2017). An intuitive explanation of this problem is that generative models have to estimate a data distribution that is usually quite complex, so the latent representations must be directly descriptive of the whole data space Loaiza-Ganem et al. (2022). This can become even more problematic for graphs because they live in a non-Euclidean and inhomogenous data space. Despite the aforementioned issues, masked autoencoding has recently shown promising results also in the graph domain with appropriately designed models (Hou et al., 2022; Tan et al., 2023).
@@
Figure 1: Illustration of the SSL approaches discussed in this paper: (a) Joint-Embedding (Contrastive) Architectures learn to create similar embeddings for inputs x and y that are compatible with each other and dissimilar embeddings otherwise. This compatibility is implemented in practice by creating different views of the input data. (b) Generative Architectures reconstruct a signal $y$ from an input signal $x$ by conditioning the decoder network on additional (potentially latent) variables $z$. (c) Joint-Embedding Predictive Architectures act as a bridge: They utilize a predictor network that processes the context $x$ and is conditioned on additional (potentially latent) variables to predict the embedding of the target $y$ in latent space.
@@
Inspired by the innovative Joint-Embedding Predictive Architecture (JEPA) (LeCun, 2022; Assran et al., 2023), we propose Graph-JEPA, a JEPA for the graph domain that can learn graph-level representations by bridging contrastive and generative models. As illustrated in Fig. 1c, a JEPA has two encoder networks that receive the input signals and produce the corresponding representations. The two encoders can potentially be different models and don’t need to share weights. A predictor module outputs a prediction of the latent representation of one signal based on the other, possibly conditioned on another variable. Graph-JEPA does not require any negative samples or complex data augmentation, and by operating in the latent space avoids the pitfalls associated with learning high-level details needed to fit the data distribution. However, the graph domain presents several challenges needed to properly design such an architecture: context and target extraction, designing a latent prediction task optimal for graph-level concepts, and learning expressive representations. In response to these questions, we equip Graph-JEPA with a specific masked modeling objective. The input graph is first divided into several subgraphs, and then the latent representation of randomly chosen target subgraphs is predicted given a context subgraph. The subgraph representations are consequently pooled to create a graph-level representation that can be used for downstream tasks.
@@
The nature of graph-level concepts is often assumed to be hierarchical (Ying et al., 2018). We conjecture that the typical latent reconstruction objective used in current JEPA formulations is not enough to provide optimal downstream performance. To this end, we design a prediction objective that starts by expressing the target subgraph encoding as a high-dimensional description of the hyperbolic angle. The predictor module is then tasked with predicting the location of the target in the 2D unit hyperbola. This prediction is compared with the target coordinates obtained by using the aforementioned hyperbolic angle. In this self-predictive setting, we explain why the stop-gradient operation and a simple predictor parametrization are useful to prevent representation collapse. To experimentally validate our approach, we evaluate Graph-JEPA against established contrastive and generative graph-level SSL methods across various graph datasets from different domains. Our proposed method demonstrates superior performance, outperforming most competitors while maintaining efficiency and ease of training. Notably, we observe from our experiments that Graph-JEPA can run up to 2.5x faster than Graph-MAE (Hou et al., 2022) and 8x faster than MVGRL (Hassani & Khasahmadi, 2020). Finally, we empirically demonstrate Graph-JEPA’s ability to learn highly expressive graph representations by showing that a linear classifier trained on the learned representations almost perfectly distinguishes pairs of non-isomorphic graphs that the 1-WL test cannot differentiate.
@@ 2.1 Self-Supervised Graph Representation Learning
Graph Neural Networks (GNNs)(Wu et al., 2019; Scarselli et al., 2008; Kipf & Welling, 2016a; Veličković et al., 2017) are now established solutions to different graph machine learning problems such as node classification, link prediction, and graph classification. Nevertheless, the cost of labeling graph data is quite high, given the immense variability of graph types and the information they can represent. To alleviate this problem, SSL on graphs has become a research frontier, where we can distinguish between two major groups of methods(Xie et al., 2022b; Liu et al., 2023):
Contrastive Methods.
@@
Contrastive learning methods usually minimize an energy function (Hinton, 2002; Gutmann & Hyvärinen, 2010) between different views of the same data. InfoGraph (Sun et al., 2019) maximizes the mutual information between the graph-level representation and the representations of substructures at different scales. GraphCL (You et al., 2020) works similarly to distance-based contrastive methods in the imaging domain. The authors first propose four types of graph augmentations and then perform contrastive learning based on them. The work of (Hassani & Khasahmadi, 2020) goes one step further by contrasting structural views of graphs. They also show that a large number of views or multiscale training does not seem to be beneficial, contrary to the image domain. Another popular research direction for contrastive methods is learning graph augmentations and how to leverage them efficiently (Suresh et al., 2021; Jin et al., 2021). Contrastive learning methods typically require a lot of memory due to data augmentation and negative samples. Graph-JEPA is much more efficient than typical formulations of these architectures, given that it does not require any augmentations or negative samples. Another major difference is that the prediction in latent space in JEPAs is done through a separate predictor network rather than using the common Siamese structure (Bromley et al., 1993)(Fig. 1a vs. 1c).
Generative Methods.
@@
The goal of generative models is to recover the data distribution, an objective that is typically implemented through a reconstruction process. In the graph domain, most generative architectures that are also used for SSL are extensions of Auto-Encoder (AE) models (Hinton & Zemel, 1993) architectures. These models learn an embedding from the input data and then use a reconstruction objective with (optional) regularization to learn the data distribution. Kipf & Welling (2016b) extended the framework of AEs and VAEs (Kingma & Welling, 2013) to graphs by using a GNN as an encoder and the reconstruction of the adjacency matrix as the training objective. However, the results on downstream tasks with embeddings learned in this way are often unsatisfactory compared with contrastive learning methods, a tendency also observed in other domains (Liu et al., 2023). A recent and promising direction is masked autoencoding (MAE) (He et al., 2022), which has proved to be a very successful framework for the image and text domains. GraphMAE (Hou et al., 2022) is an instantiation of MAEs in the graph domain, where the node attributes are perturbed and then reconstructed, providing a paradigm shift from the structure learning objective of GAEs. S2GAE (Tan et al., 2023) is one of the latest GAEs, which focuses on reconstructing the topological structure but adds several auxiliary objectives and additional designs. Our architecture differs from generative models in that it learns to predict directly in the latent space, thereby bypassing the necessity of remembering and overfitting high-level details that help maximize the data evidence (Fig. 1b vs. 1c).
@@ 2.2 Joint-Embedding Predictive Architectures
Joint-Embedding Predictive Architectures (LeCun, 2022) are a recently proposed design for SSL. The idea is similar to both generative and contrastive approaches, yet JEPAs are non-generative since they cannot directly predict $y$ from $x$, as shown in Fig. 1c. The energy of a JEPA is given by the prediction error in the embedding space, not the input space. These models can intuitively be understood as a way to capture abstract dependencies between $x$ and $y$, potentially given another latent variable $z$. It is worth noting that the different models comprising the architecture may differ in terms of structure and parameters. An in-depth explanation of Joint-Embedding Predictive Architectures and their connections to human representation learning is provided by LeCun (2022). Some works acknowledged that latent self-predictive architectures were effective (Grill et al., 2020; Chen & He, 2021) even before JEPAs effectively became synonymous with this concept. Inspired by these trends, a number of related works have tried to employ latent prediction objectives for graph SSL, showing advantages mostly compared to contrastive learning. Thakoor et al. (2021) perform latent self-prediction on augmented views of a graph in a similar fashion to BYOL (Grill et al., 2020), while Zhang et al. (2021) rely on ideas from Canonical Correlation Analysis to frame a learning objective that preserves feature invariance and forces decorrelation when necessary. The work of Lee et al. (2022) presents a model that learns latent positive examples through a k-NN and clustering procedure in the transductive setting, while Xie et al. (2022a) combine instance-level reconstruction (generative pretraining) and feature-level invariance (latent prediction). Given that these models learn using a latent self-predictive objective, similar to ours, we will refer to them also using the term self-predictive in the rest of the paper. Unlike previously proposed methods, Graph-JEPA operates exclusively in the latent space and implements a novel training task without the need for data augmentation. At the current state of the art, the JEPA framework has been implemented for images (Assran et al., 2023), video(Bardes et al., 2023b; a), and audio (Fei et al., 2023). We propose the first architecture implementing the modern JEPA principles for the graph domain and use it to learn graph-level representations.
@@ 3 Method
A general overview.
We consider graphs $G$ defined as $G=(V,E)$ where $V=\{v_{1}\ldots v_{N}\}$ is the set of nodes, with a cardinality $|V|=N$, and $E=\{e_{1}\ldots e_{M}\}$ is the set of edges, with a cardinality $|E|=M$. For simplicity of the exposition, we consider symmetric, unweighted graphs, although our method can be generalized to weighted or directed graphs. In this setting, $G$ can be represented by an adjacency matrix $A\in\{0,1\}^{N\times N}$, with $A_{ij}=1$ if nodes $v_{i}$ and $v_{j}$ are connected and 00 otherwise. Fig. 2 provides the overview of the proposed architecture. The high-level idea of Graph-JEPA is to divide the input graph into subgraphs (patches) (He et al., 2023) and then predict the representation of a randomly chosen target subgraph from the representation of a single context subgraph (Assran et al., 2023). Again, we would like to stress that this masked modeling objective is realized in latent space without needing negative (or positive) samples. The subgraph representations are then pooled to create a vector representation for the whole graph, i.e., a graph-level representation. Therefore, the learning procedure consists of a sequence of operations: 1. Spatial Partitioning; 2. Subgraph Embedding; 3. Context and Target Encoding; 4. Latent Target Prediction.
Figure 2: A complete overview of Graph-JEPA. We first extract non-overlapping subgraphs (patches) (a.), perform a 1-hop neighborhood expansion (b.), and encode the subgraphs with a GNN (c.). After the subgraph encoding, one is randomly picked as the context and $m$ others as the targets (d.), and they are fed into their respective encoders (e.). The embeddings generated from the target encoder are used to produce the target subgraphs’ coordinates $\psi_{y}$. Finally, the predictor network is tasked with directly predicting the coordinates $\hat{\psi}_{y}$ for each target subgraph based on the context embedding and the positional embedding of each target subgraph (f.). A regression loss acts as the energy function $D$ between the predicted and target coordinates. Note that the extracted subgraphs in (a.) and (b.) are meant for illustrative purposes only. The number of nodes in each subgraph can vary.